Automatic Management System for the AtlasCar Gearbox

Automatic Management System for the AtlasCar Gearbox

Pedro Filipe Pinto Pinheiro

Department of Mechanical Engineering University of Aveiro

July 9, 2012

Automatic Management System for the AtlasCar Gearbox

Outline

1 Introduction

The Atlas Project

The AtlasCar Actuators The Problem

2 Objectives
3 DC Motor Controller
4 Firmware Programming
5 Hardware-in-the-Loop Simulator

Automatic Management System for the AtlasCar Gearbox

Introduction The Atlas Project

Outline

1 Introduction

The Atlas Project

The AtlasCar Actuators The Problem

2 Objectives
3 DC Motor Controller
4 Firmware Programming
5 Hardware-in-the-Loop Simulator

The Atlas Project

ADAS;

Safety Systems; Data Gathering; Autonomous Navigation.

Automatic Management System for the AtlasCar Gearbox

Introduction The AtlasCar Actuators

Outline

1 Introduction

The Atlas Project

The AtlasCar Actuators The Problem

2 Objectives
3 DC Motor Controller
4 Firmware Programming
5 Hardware-in-the-Loop Simulator

Automatic Management System for the AtlasCar Gearbox

Introduction The AtlasCar Actuators

The AtlasCar Actuators

Throttle; Brake pedal; Clutch pedal; Handbrake; Ignition; Steering; Lights.

Automatic Management System for the AtlasCar Gearbox

Introduction The Problem

Outline

1 Introduction

The Atlas Project

The AtlasCar Actuators The Problem

2 Objectives
3 DC Motor Controller
4 Firmware Programming
5 Hardware-in-the-Loop Simulator

The Problem

What about the gearbox?

Automatic Management System for the AtlasCar Gearbox Objectives

Objectives

Slight changes on the Mechanical System; Development of a Control PCB; Controller Firmware; Robust Communication Messaging System; System Simulation and Laboratory Tests.

Automatic Management System for the AtlasCar Gearbox DC Motor Controller

Arduino DC Motor Controller

The controller must: DC Motors Power Controller; Read both multi-turn potentiometers; Read the MAN/AUTO switches; Control Seven Segment Display.

The Power Controller

H-bridge IC VNH3SP30-E;

▶ 12 V; ▶ 20 A.

Coded Inputs.